# General targeted machine learning for modern causal mediation analysis

### v1

@@ Abstract
Causal mediation analyses investigate the mechanisms through which causes exert their effects, and are therefore central to scientific progress. The literature on the non-parametric definition and identification of mediational effects in rigourous causal models has grown significantly in recent years, and there has been important progress to address challenges in the interpretation and identification of such effects. Despite great progress in the causal inference front, statistical methodology for non-parametric estimation has lagged behind, with few or no methods available for tackling non-parametric estimation in the presence of multiple, continuous, or high-dimensional mediators. In this paper we show that the identification formulas for six popular non-parametric approaches to mediation analysis proposed in recent years (natural direct and indirect effects, randomized interventional effects, separable effects, organic direct and indirect effects, recanting twin effects, and decision theoretic effects) can be recovered from just two statistical estimands. We leverage this finding to propose an all-purpose one-step estimation algorithm that can be coupled with machine learning in any mediation study that uses any of these six definitions of mediation. The estimators rely on a re-parameterization of the identification formulas in terms of sequential regressions, and on a first order non-parametric von-Mises approximations of the first bias of a plug-in estimator. This allows us to show that the proposed one-step estimators have desirable properties, such as $\sqrt{n}$-convergence and asymptotic normality. Estimating the first-order correction for the one-step estimator requires estimation of complex density ratios on the potentially high-dimensional mediators, a challenge that is solved using recent advancements in so-called Riesz learning. We illustrate the properties of our methods in a simulation study and illustrate its use on real data to estimate the extent to which pain management practices mediate the total effect of having a chronic pain disorder on opioid use disorder. We provide an R package implementing our methods publicly available at https://github.com/nt-williams/crumble.
@@ 1 Introduction
Natural direct and indirect effects (NDE and NIE; Robins and Greenland, 1992; Pearl, 2001) are one of the most widely known definitions of causal mediation parameters with an agreed upon mechanistic interpretation. Broadly, the NDE measures the effect that operates independent of a mediator through counterfactual variables in which the effect through the mediator is disabled, whereas the NIE measures effects through the mediator by considering counterfactuals that disable all effects that operate independently of it. In spite of their scientific importance, desirable interpretation, and widespread use, the NDE and NIE are seldom identified from observed data, because their identification relies on so-called cross-world counterfactual assumptions—independence assumptions between counterfactual variables indexed by distinct interventions on the exposure. Philosophically, cross-world assumptions are problematic because they cannot be tested empirically nor enforced by design, which means that conclusions from an analysis using the NIE and NDE are not falsifiable. Practically, the cross-world assumption is guaranteed not to hold in the presence of variables caused by the treatment that are common causes of both the mediator and the outcome (Avin et al., 2005), regardless of whether such variables are measured or not. These variables, often referred to as intermediate confounders or recanting witnesses, are pervasive in scientific applications.
@@
In an attempt to address the above limitations of the NDE and NIE, several alternative causal estimands have been proposed. In this article we discuss five such definitions, but we note that our review is not exhaustive. First, randomized interventional direct and indirect effects (RIDE and RIIE van der Laan and Petersen, 2008; VanderWeele et al., 2014; Vansteelandt and Daniel, 2017; VanderWeele and Tchetgen Tchetgen, 2017; Díaz et al., 2023) define direct and indirect effects through interventions that set the mediator to a random draw based on the distributions of the counterfactual mediators under treatment and control. Randomized interventional effects are identifiable in the presence of intermediate confounders under the assumption that all confounders of the exposure-outcome, exposure-mediator, and mediator-outcome relations are measured. However, it was recently shown that randomized interventional effects may not necessarily measure mechanistic relations, as the RIIE could be nonzero even in a situation where some population units experience an exposure-mediator effect, other units experience a mediator-outcome effect, but no unit experiences both (Miles, 2023). Second, in a series of articles, Didelez et al. (2006); Geneletti (2007) and Dawid (2021) proposed a decision-theoretic approach to mediation analysis (and more generally to causal inference) that uses probabilistic definitions for the relevant interventions, avoiding reliance on on counterfactual arguments, and therefore avoiding reliance on cross-world independence assumptions. Third, a recent strand of the causal inference literature has addressed the definition and identification of mediational effects and effects under competing events using so-called separable effects (Robins and Richardson, 2010; Didelez, 2019; Stensrud et al., 2021; Robins et al., 2022; Stensrud et al., 2022). These effects avoid cross-world definitions by requiring researchers to posit the existence of separable components of the exposure that exert independent effects on the outcome and the mediator. Separable effects are then defined through distinct interventions on those components, thus avoiding cross-world statements. Vo et al. (2024) generalized separable effects to allow for the definition and identification of mediational effects in the presence of intermediate confounders. Fourth, Díaz (2024) and Vo et al. (2024) recently introduced a four-way decomposition of the average treatment effect into each of the path-specific effects involved, using randomized versions of the recanting witnesses. Unlike some of the previous approaches, these path-specific effects satisfy a property known as the path-specific sharp null criterion (Miles, 2023), which guarantees that the effects are zero whenever no individual in the population experiences an effect through the corresponding path. Lastly, a series of articles discusses organic direct and indirect effects (Lok, 2015, 2016, 2019; Lok and Bosch, 2021), which rely on the characterization of an external intervention to the causal system that would yield specific modifications to the mediator distribution.
An additional limitation of all of the above approaches for the definition, identification, and estimation of direct and indirect effects in mediation analyses is that the literature has almost exclusively focused on binary exposures, although some approaches for continuous and/or multivariate exposures have been developed based on stochastic interventions and modified treatment policies (Díaz and Hejazi, 2020; Hejazi et al., 2023; Gilbert et al., 2024).
As illustrated above, the non-parametric definition and identification of causal effects for mediation analysis has received significant attention which has led to significant progress. Some methods for the non-parametric estimation of those effects have also been developed. For instance, non-parametric approaches to estimate the statistical functional identifying the NDE and NIE include the estimators of van der Laan and Petersen (2008); Zheng and van der Laan (2012), and Xia and Chan (2023). Approaches for randomized interventional effects for cross-sectional (Díaz et al., 2021a; Benkeser and Ran, 2021; Rudolph et al., 2024a) and longitudinal (Zheng and van der Laan, 2017; Díaz et al., 2023; Wang et al., 2023) effects have also been developed, as well as non-parametric estimators for the NDE and NIE under alternative monotonicity identifying assumptions that do not preclude intermediate confounders (Rudolph et al., 2023; Tchetgen and VanderWeele, 2014).
In spite of important progress in the non-parametric definition and identification of mediational parameters, estimators for many approaches remain underdeveloped, and there are few solutions that can handle continuous and multivariate or high-dimensional mediators. Specifically, most of the above literature on non-parametric estimation of mediational parameters is based on recently popularized frameworks that estimate and/or correct for the first-order bias of a plug-in machine learning estimator (van der Laan and Rose, 2011, 2018; Chernozhukov et al., 2018). Central to these frameworks is the concept of a canonical gradient or efficient influence function, which characterizes the first-order bias of plug-in estimators as well as the non-parametric efficiency bound (von Mises, 1947; Pfanzagl and Wefelmeyer, 1985; Bickel et al., 1997). A key challenge in mediation analysis with continuous or high-dimensional mediators is that estimation of the canonical gradient requires estimation of densities of the mediator as well as integrals with respect to such densities. Of the estimators mentioned in the above paragraph that accommodate intermediate confounders, only the approach of Rudolph et al. (2024a) for randomized interventional effects allows for high-dimensional or multivariate mediators and confounders. As noted earlier, however, randomized interventional effects do not always allow for an interpretation in terms of individual-level mechanisms. Thus, there is a need in the literature to develop estimators of alternative approaches that address the shortcomings of randomized interventional effects, such as the separable effects Vo et al. (2024) and the recanting twin effects of Díaz (2024).
In this article we make two main contributions. First, we show that the identification formulas for the six causal parameters for mediation analysis with binary exposures described above and listed in the abstract can be recovered from two fundamental statistical functionals. We then use this fact to propose unified theory for semiparametric estimation of mediation analyses, providing an algorithm that can be used under either of these six causal models and definitions, without restrictions on the dimensionality of the mediators or the intermediate confounders. We address challenges in the estimation of multivariate or high-dimensional densities on the mediator through the use of so-called Riesz learning (Chernozhukov et al., 2022a), and address integration with respect to those densities through parameterizing the integrals as conditional expectations using stratified random permutations of the observed variables as in Doran et al. (2014). Second, we provide an extension of these methods to definitions of mediational parameters for continuous or multivariate exposures, by means of modified treatment policies (Haneuse and Rotnitzky, 2013; Sani et al., 2020; Díaz et al., 2021b, 2022).
@@
The paper is organized as follows. In §2, we introduce general identification formulas for mediation analysis under a binary exposure, which unifies all six causal parameters mentioned before. In §3, we introduce the semi-parametric efficiency theory for such parameters. In §4, we use the developed theory to introduce estimation strategies. We prove desirable properties of the estimators such as asymptotic linearity, weak convergence, and double robustness. In §5, we demonstrate how the methods can be extended to non-binary treatments using modified treatment policies. In §6, we present numerical illustrations of the performance of the estimators in simulated datasets. An illustrative real data example investigating the effects of chronic pain disorder on opioid use disorder is given in §7.
@@ 2 General identification formulas for mediation analysis with binary exposures
For values $(a_{1},\ldots,a_{4})\in\{0,1\}^{4}$, we define
$\psi^{R}(a_{1},a_{2},a_{3},a_{4})$ $=\mathsf{E}\Bigg{[}\mathsf{E}\bigg{[}\mathsf{E}\Big{[}\mathsf{E}\big{[}\mathsf{E}(Y\,|\,A=a_{1},Z^{\pi},M,W)\,\big{|}\,A=a_{2},M,W\big{]}\,\Big{|}\,A=a_{3},Z,W\Big{]}\,\bigg{|}\,A=a_{4},W\bigg{]}\Bigg{]},$ (2)
$\psi^{N}(a_{1},a_{2},a_{3})$ $=\mathsf{E}\bigg{[}\mathsf{E}\Big{[}\mathsf{E}\big{[}\mathsf{E}(Y\mid A=a_{1},Z,M,W)\,\big{|}\,A=a_{2},Z,W\big{]}\,\Big{|}\,A=a_{3},W\Big{]}\bigg{]},$
where the superscript $N$ stands for natural and the superscript $R$ stands for randomized, accounting for the fact that we introduce the permutation $Z^{\pi}$. The introduction of the permuted variable $Z^{\pi}$ allows us to write integrals of the type $\int f(a_{1},z,m,w)\,\mathrm{d}\mathsf{P}(z\mid a_{2},w)$, which may be hard to estimate if $Z$ is high-dimensional, as conditional expectations $\mathsf{E}[f(a_{1},Z^{\pi},M,W)\mid A=a_{2},M=m,W=w]$, which can be estimated using flexible regression techniques from the statistical and machine learning literature.
In Proposition 1 below we show that the identifying functionals for the six mediation frameworks described in the abstract can be recovered from the parameters $\psi^{R}$ and $\psi^{N}$ defined above. In Table 1 below we introduce abbreviations for each of the corresponding causal effects, and provide one reference for each where the corresponding definitions in terms of causal models and proofs for identification can be found. Of the approaches listed in Table 1, only the first two do not allow for intermediate confounders (i.e., these two approaches assume there is no variable $Z$ in the DAG in Figure 1), and only the last two allow for a decomposition of the average treatment effect into path-specific effects. The parameters referred to as “indirect effects” measure all effects of $A$ on $Y$ that operate through $M$ (i.e., $A\to M\to Y$ and $A\to Z\to M\to Y$), and the parameters referred to as “direct effects” measure all effects of $A$ on $Y$ that operate independently of $M$ (i.e., $A\to Z\to Y$ and $A\to Y$). For approaches that allow a decomposition into path-specific effects, we introduce the notation $P_{j}$ to refer to each path, where $P_{1}:A\rightarrow Y$, $P_{2}:A\rightarrow Z\rightarrow Y$, $P_{3}:A\rightarrow Z\rightarrow M\rightarrow Y$, and $P_{4}:A\rightarrow M\rightarrow Y$. A comprehensive discussion of the causal models, definitions of mediational effects, and identification assumptions, is out of the scope of this paper and is given in the references provided in §1 and in Table 1, we refer the readers to those references for discussion.
@@
Having established the importance and generality of the parameters $\psi^{N}$ and $\psi^{R}$ for mediation analysis, we will proceed to develop general estimation methods for those parameters that can be used with high-dimensional $W$, $Z$, and $M$ variables. In what follows, we ease notation by omitting $(a_{1},a_{2},a_{3},a_{4})$ when referring to $\psi^{N}$ and $\psi^{R}$, but the dependence of the parameters on these values will remain implicit. In order to develop non-parametric estimators that can leverage machine learning to fit the regressions involved in the definitions of $\psi^{N}$ and $\psi^{R}$, we will use techniques from the literature in semi-parametric efficiency theory, which require the study of certain properties of the parameters $\psi^{N}$ and $\psi^{R}$ seen as functionals $\psi^{N}(\mathsf{P})$ and $\psi^{R}(\mathsf{P})$ that map a probability distribution $\mathsf{P}$ in the non-parametric model $\cal P$ into a real number given by the formulas in (2). The following section provides a study of the required elements.

### v2

@@ Abstract
Causal mediation analyses investigate the mechanisms through which causes exert their effects, and are therefore central to scientific progress. The literature on the non-parametric definition and identification of mediational effects in rigorous causal models has grown significantly in recent years, and there has been important progress to address challenges in the interpretation and identification of such effects. However, statistical methodology for non-parametric estimation has lagged, with few or no methods available for tackling non-parametric estimation in the presence of multiple, continuous, or high-dimensional mediators. In this paper we propose an all-purpose one-step estimation algorithm that can be coupled with machine learning in mediation studies that use any of six common mediation parameters (natural direct and indirect effects, randomized interventional effects, separable effects, organic direct and indirect effects, recanting twin effects, and decision theoretic effects). The estimators build on methods for double machine learning, including a re-parameterization of the identification formulas in terms of sequential regressions, a first order non-parametric von-Mises approximation of the first-order bias of a plug-in estimator, and Riesz learning for estimation of nuisance parameters. We show that the proposed one-step estimators have desirable properties, such as $\sqrt{n}$-convergence and asymptotic normality. We illustrate the properties of our methods in a simulation study and demonstrate its use on real data to estimate the extent to which pain management practices mediate the total effect of having a chronic pain disorder on opioid use disorder. We provide an R package (on CRAN) implementing our methods publicly available at https://github.com/nt-williams/crumble.
@@ 1.1 Prior literature
Natural direct and indirect effects (NDE and NIE; Robins and Greenland, 1992; Pearl, 2001) are one of the most widely known definitions of causal mediation parameters with an agreed upon mechanistic interpretation. Broadly, the NDE measures the effect that operates independent of a mediator through counterfactual variables in which the effect through the mediator is disabled, whereas the NIE measures effects through the mediator by considering counterfactuals that disable all effects that operate independently of it. In spite of their scientific importance, desirable interpretation, and widespread use, the NDE and NIE are seldom identified from observed data, because their identification relies on so-called cross-world counterfactual assumptions—independence assumptions between counterfactual variables indexed by distinct interventions on the exposure. Philosophically, cross-world assumptions are problematic because they cannot be tested empirically nor enforced by design, which means that conclusions from an analysis using the NIE and NDE are not falsifiable. Practically, the cross-world assumption is guaranteed not to hold in the presence of variables caused by the treatment that are common causes of both the mediator and the outcome (Avin et al., 2005), regardless of whether such variables are measured or not. These variables, often referred to as intermediate confounders, are pervasive in scientific applications. Although Vo et al. (2024) proposed a test to confirm whether a given variable $Z$ is an intermediate confounder, this test cannot rule out the existence of unmeasured intermediate confounders that would invalidate the cross-world assumption.
@@
In an attempt to address the above limitations of the NDE and NIE, several alternative causal parameters have been proposed. In this article we discuss five such definitions, but we note that our review is not exhaustive. First, randomized interventional direct and indirect effects (RIDE and RIIE; van der Laan and Petersen, 2008; VanderWeele et al., 2014; Vansteelandt and Daniel, 2017; VanderWeele and Tchetgen Tchetgen, 2017; Díaz et al., 2023) define direct and indirect effects through interventions that set the mediator to a random draw based on the distributions of the counterfactual mediators under treatment and control. Randomized interventional effects are identifiable in the presence of intermediate confounders under the assumption that all confounders of the exposure-outcome, exposure-mediator, and mediator-outcome relations are measured. However, it was recently shown that randomized interventional effects may not necessarily measure mechanistic relations, as the RIIE could be nonzero even in a situation where some population units experience an exposure-mediator effect, other units experience a mediator-outcome effect, but no unit experiences both (Miles, 2023). Second, in a series of articles, Didelez et al. (2006); Geneletti (2007) and Dawid (2021) proposed a decision-theoretic approach to mediation analysis (and more generally to causal inference) that uses probabilistic definitions for the relevant interventions, avoiding reliance on counterfactual arguments, and therefore avoiding reliance on cross-world independence assumptions. Third, a recent strand of the causal inference literature has addressed the definition and identification of path-specific effects and effects under competing events using so-called separable effects (Robins and Richardson, 2010; Didelez, 2019; Stensrud et al., 2021; Robins et al., 2022; Stensrud et al., 2022). These effects avoid cross-world definitions by requiring researchers to posit the existence of separable components of the exposure that exert independent effects on the outcome and the mediator. Separable effects are then defined through distinct interventions on those components, thus avoiding cross-world statements. Vo et al. (2024) generalized separable effects to allow for the definition and identification of mediational effects in the presence of intermediate confounders. Fourth, Díaz (2024) and Vo et al. (2024) recently introduced a four-way decomposition of the average treatment effect into each of the path-specific effects involved, using randomized versions of the recanting witnesses. Unlike some of the previous approaches, these path-specific effects satisfy a property known as the path-specific sharp null criterion (Miles, 2023), which guarantees that the effects are zero whenever no individual in the population experiences an effect through the corresponding path. Lastly, a series of articles discusses organic direct and indirect effects (Lok, 2015, 2016, 2019; Lok and Bosch, 2021), which rely on the characterization of an external intervention to the causal system that would yield specific modifications to the mediator distribution.
1.2 Our contribution
In spite of important progress in the non-parametric definition and identification of mediational parameters, estimators for many approaches remain underdeveloped, and there are few non-parametric solutions [^1] See van der Laan and Petersen (2008); Zheng and van der Laan (2012); Xia and Chan (2023) for estimators of the NDE/NIE. See Díaz et al. (2021a); Benkeser and Ran (2021); Rudolph et al. (2024a) and Zheng and van der Laan (2017); Díaz et al. (2023); Wang et al. (2023) for randomized interventional effects for cross-sectional and longitudinal settings, respectively. with only a handful that can handle continuous or high-dimensional mediators in specific settings. [^2] To our knowledge, only the approach of Rudolph et al. (2024a) for randomized interventional effects allows for high-dimensional mediators and intermediate confounders. Methods for mediation analyses that can handle continuous or high-dimensional mediators are paramount in contemporary research fields such as the study of so-called omics, (e.g., the microbiome is often considered a mediator of effects on health (Sohn and Li, 2019; Wang et al., 2020)), studies of fairness in machine learning (Nabi and Benkeser, 2024), and neuroscience (Chén et al., 2018), among many others. An additional limitation of all of the above approaches for the definition, identification, and estimation of direct and indirect effects in mediation analyses is that the literature has almost exclusively focused on binary exposures, although some approaches for continuous and/or multivariate exposures have been developed based on stochastic interventions and modified treatment policies (Díaz and Hejazi, 2020; Hejazi et al., 2023; Gilbert et al., 2024). In contrast, we resolved the estimation issue by making two main contributions. First, we developed non-parametric double machine learning estimators for all six mediation parameters listed in the abstract for the case of binary treatments in the presence of continuous or multivariate, mediators and intermediate confounders. Second, we proposed two approaches (based on randomized interventional effects, recanting-twins effects) for defining mediation parameters in the presence of continuous or multivariate treatments, and proposed estimation methods that can handle continuous or multivariate, mediators and intermediate confounders.
Our estimators are based on recently popularized frameworks that estimate and/or correct for the first-order bias of a plug-in machine learning estimator (van der Laan and Rose, 2011, 2018; Chernozhukov et al., 2018). Central to these frameworks is the concept of a canonical gradient or efficient influence function (EIF), which characterizes the first-order bias of plug-in estimators as well as the non-parametric efficiency bound (von Mises, 1947; Pfanzagl and Wefelmeyer, 1985; Bickel et al., 1997). A key challenge in mediation analysis with continuous or high-dimensional mediators is that estimation of the canonical gradient requires estimation of densities of the mediator as well as integrals with respect to such densities. We address challenges in the estimation of multivariate or high-dimensional densities on the mediator through the use of so-called Riesz learning (chernuzhukov2022riesznet), and address integration with respect to those densities through parameterizing the integrals as conditional expectations using stratified random permutations of the observed variables as in Doran et al. (2014). Definitions of mediation parameters for continuous or multivariate exposures leverage modified treatment policies (Haneuse and Rotnitzky, 2013; Sani et al., 2020; Díaz et al., 2021b, 2022), also called dynamic interventions that depend on the natural value of treatment (Robins et al., 2004; Richardson and Robins, 2013; Young et al., 2014).
@@ 1.2 Our contribution
The rest of the paper is organized as follows. In §2, we introduce general identification formulas for mediation analysis under a binary exposure, where we use the fact that the identification formulas for the six causal parameters for mediation analysis with binary exposures listed in the abstract can be recovered from two fundamental statistical functionals. In §3, we introduce the semi-parametric efficiency theory these two functionals. In §4, we use the developed theory to introduce estimation strategies. We prove desirable properties of the estimators such as asymptotic linearity, weak convergence, and double robustness. In §5 we present definitions for non-binary treatments using modified treatment policies and introduce the corresponding estimators. In §6, we present numerical illustrations of the performance of the proposed approach in simulated datasets. An illustrative real data example investigating the effects of chronic pain disorder on opioid use disorder is given in §7. A discussion about this manuscript and future work is given in §8.
@@ 2 General identification formulas for mediation analysis with binary exposures
Let $\theta(A,Z,M,W)=\mathsf{E}[Y\mid A,Z,M,W]$. For values $(a_{1},\ldots,a_{4})\in\{0,1\}^{4}$, we define
$\psi^{R}(a_{1},a_{2},a_{3},a_{4})$ $=\mathsf{E}\Bigg{[}\mathsf{E}\bigg{[}\mathsf{E}\Big{[}\mathsf{E}\big{[}\theta(a_{1},Z^{\pi},M,W)\,\big{|}\,A=a_{2},M,W\big{]}\,\Big{|}\,A=a_{3},Z,W\Big{]}\,\bigg{|}\,A=a_{4},W\bigg{]}\Bigg{]},$ (2)
$\psi^{N}(a_{1},a_{2},a_{3})$ $=\mathsf{E}\bigg{[}\mathsf{E}\Big{[}\mathsf{E}\big{[}\theta(a_{1},Z,M,W)\,\big{|}\,A=a_{2},Z,W\big{]}\,\Big{|}\,A=a_{3},W\Big{]}\bigg{]},$
where the superscript $N$ stands for natural and the superscript $R$ stands for randomized, accounting for the fact that we introduce the permutation $Z^{\pi}$. In Proposition 1 below we show that the identifying functionals for the six mediation frameworks described in the abstract can be recovered from just two parameters $\psi^{R}$ and $\psi^{N}$ defined above. The introduction of the permuted variable $Z^{\pi}$ allows us to write integrals of the type $\int f(a_{1},z,m,w)\,\mathrm{d}\mathsf{P}(z\mid a_{2},w)$, which may be hard to estimate numerically if $Z$ is continuous or high-dimensional, as conditional expectations $\mathsf{E}[f(a_{1},Z^{\pi},M,W)\mid A=a_{2},M=m,W=w]$, which can be estimated using flexible regression techniques from the statistical and machine learning literature.
In Table 1 below we introduce abbreviations for each of the corresponding causal effects, and provide one reference for each where the corresponding definitions in terms of causal models and proofs for identification can be found. Of the approaches listed in Table 1, only the first two do not allow for intermediate confounders (i.e., these two approaches assume there is no variable $Z$ in the DAG in Figure 1), and only the last two allow for a decomposition of the average treatment effect into path-specific effects. The parameters referred to as “indirect effects” measure all effects of $A$ on $Y$ that operate through $M$ (i.e., $A\to M\to Y$ and $A\to Z\to M\to Y$), and the parameters referred to as “direct effects” measure all effects of $A$ on $Y$ that operate independently of $M$ (i.e., $A\to Z\to Y$ and $A\to Y$). For approaches that allow a decomposition into path-specific effects, we introduce the notation $P_{j}$ to refer to each path, where $P_{1}:A\rightarrow Y$, $P_{2}:A\rightarrow Z\rightarrow Y$, $P_{3}:A\rightarrow Z\rightarrow M\rightarrow Y$, and $P_{4}:A\rightarrow M\rightarrow Y$. A comprehensive discussion of the causal models, definitions of mediational effects, and identification assumptions, is out of the scope of this paper and is given in the references provided in §1 and in Table 1, we refer the readers to those references for discussion.
@@
Remark 1.
The recanting twins effects discussed in Vo et al. (2024) are generally different from the natural path-specific effects. However, recanting twin effects have desirable properties that recover features of natural path-specific effects, such as 1) monotonicity, meaning they increase or decrease in alignment with the natural path-specific effects, 2) have the same sign with the natural path-specific effects, and 3) recover natural path-specific effects in the absence of intermediate confounding.
A proof of Proposition 1 is shown in the supplementary materials. Having established the importance and generality of the parameters $\psi^{N}$ and $\psi^{R}$ for mediation analysis, we will proceed to develop general estimation methods for those parameters that can be used with high-dimensional $W$, $Z$, and $M$ variables. In what follows, we ease notation by omitting $(a_{1},a_{2},a_{3},a_{4})$ when referring to $\psi^{N}$ and $\psi^{R}$, but the dependence of the parameters on these values will remain implicit. In order to develop non-parametric estimators that can leverage machine learning to fit the regressions involved in the definitions of $\psi^{N}$ and $\psi^{R}$, we will use techniques from the literature in semi-parametric efficiency theory, which require the study of certain properties of the parameters $\psi^{N}$ and $\psi^{R}$ seen as functionals $\psi^{N}(\mathsf{P})$ and $\psi^{R}(\mathsf{P})$ that map a probability distribution $\mathsf{P}$ in the non-parametric model $\cal P$ into a real number given by the formulas in (2). The following section provides a study of the required elements.